itzsummer:

During a science experiment juanita recorded that she blinked 11 times in one minute if this is a normal count and juanita wakes up at 7 A.M. and goes to bed at 10 P.M. how many times will she blink during the time she is awake?

itzsummer:

A 165 B 660 C 9900 D 15,840

photonics:

\(\large 11 \frac{times}{min}*60\frac{min}{hr}*(22-7) hrs=9900 \text{ blinks}\)

Shadow:

@itzsummer Do you understand how photonics solved this problem?

itzsummer:

yes thanks so much

Shadow:

If you wish, you can medal him. He barely answers any questions, so I was surprised that he answered this one. @itzsummer

itzsummer:

how do i do that

itzsummer:

how its that sign beside the star

itzsummer:

he's really smart wow lol

Shadow:

Yes he is.